Determine the [H+] , pH, and pOH of a solution with an [OH−] of 1.9×10−13M at 25 °C.

Determine the [H+] , [OH−] , and pOH of a solution with a pH of 4.77 at 25 °C.

Determine the [H+] , [OH−] , and pH of a solution with a pOH of 11.38 at 25 °C.
1
Expert's answer
2020-04-15T01:52:24-0400
  1. pOH=-lg[OH-]= 12.72

pH=14-pOH=1.28

[H+]=10^-pH=0.05248

2.pOH=9.23

[H+]=1.698*10^-5

[OH-]=5.88*10^-10

3.pH=2.62

[H+]=0.0024=2.4*10^-3

[OH-]=4.17*10^-12